OpenStudy (anonymous):

The width of a rectangle is 7cm shorter than its length and its area is 228 sq. cm. Find the dimensions f the rectangle.

OpenStudy (anonymous):

x(x-7) =228 ==> x^2-7x-228 =0 ===> (x+12)(x-19)=0 ===> x=-12 or x=19 ===> dimentions are 19 and (19-7 =12)